# IEEEexample:BSTcontrol

### v1

@@ Abstract
This paper investigates the capability of Neural Networks (NNs) as alternatives to the traditional methods to analyse the performance of aerofoils used in the wind and tidal energy industry. The current methods used to assess the characteristic lift and drag coefficients include Computational Fluid Dynamics (CFD), thin aerofoil and panel methods, all face trade-offs between computational speed and the accuracy of the results and as such NNs have been investigated as an alternative with the aim that it would perform both quickly and accurately. As such, this paper provides a benchmark for the windAI_bench dataset [1] published by the National Renewable Energy Laboratory (NREL) in the USA. In order to validate the methodology of the benchmarking, the AirfRANS dataset [2] is used as both a starting point and a point of comparison. This study evaluates four neural networks (MLP, PointNet, GraphSAGE, GUNet) trained on a range aerofoils at 25 angles of attack (4 to 20). to predict fluid flow and calculate lift coefficients ($C_{L}$) via the panel method. GraphSAGE and GUNet performed well during the testing phase, but underperformed during validation. Accordingly, this paper has identified PointNet and MLP as the two strongest models tested, however whilst the results from MLP are more commonly correct for predicting the behaviour of the fluid, the results from PointNet provide the more accurate results for calculating $C_{L}$.
@@ 1 Introduction
In order to validate the methodology applied to benchmarking windAI_bench, the second dataset being investigated is the AirfRANS dataset published by Bonnet et al. [2]. AirfRANS contains 1,000 aerofoils and is benchmarked for use in Machine Learning (ML) by using four separate architectures. Importantly, whilst it includes training and testing loops, AirfRANS does not incorporate validation with unseen shapes and outside the aerofoils shape distribution. This means that the ability of any of the benchmark models to generalise to completely unseen data has not been verified. To keep the results in line with those for similar datasets, the models used will be the same as those in Bonnet et al. [2]. One point worth noting is that whilst these datasets are designed for wind turbines, they are also applicable to the design and evaluation of tidal turbines. To account for scalability issues, Reynolds number calculations would thus need to consider flow characteristics and operational velocities at which tidal turbines usually operate.
As mentioned, the computational requirement of generating CFD data means that the data is regularly unattainable in large enough quantities for data-driven NNs to be viable. This paper aims to investigate the amount of data required to attain reasonable results. Therefore, the performance of each model has been investigated over 5, 20, 55 and 150 aerofoils. These numbers were chosen to be roughly equal to an exponential scale as $5\approx e^{1.5}$, $20\approx e^{3}$, $55\approx e^{4}$ and $150\approx e^{5}$. It is important to validate the results from this against a sample set of aerofoils used neither in the training nor testing of the model, i.e., a validation dataset, to ensure that the model is not over-fitting to the data presented.
@@ 2 Related Literature
Artificial Neural Networks (ANNs), or simply NNs, are universal function approximators [10] used within ML for the prediction of specified functions. NNs are called so because they are based on the design of the biological neural networks found within a brain, with neurons and synapses represented by the neurons that make up each layer of the NN and the connections that connect them, a diagram of which can be seen in Figure 1. The network is divided up into three categories of layers (input, hidden and output) each of a user specified width, i.e. how many neurons in each layer. Each connection contains a weight ($w$) which can be within the range $0\leq w\leq 1$. These are multiplied by the output from the neuron it comes from ($x$) before being summed up with the remaining connections along with a bias factor ($b$), as seen in Eq. 1.
@@
These weights and neuron values are used to affect the final output of the NN and this output is then compared to a ‘ground truth’ value, i.e. the ‘real’ value usually taken from the data. Then, after finding the gradient of the ‘Loss’, or difference between these two values, the weights and biases of the network are updated using an optimisation algorithm such as ADAM [11]. The data loaded into the model is split into training and testing data: the former is used to optimise the weights and biases of the model: whilst the latter is used to check the ability of the model to generalise to unseen data. It is worth noting that incorporating a testing split of data is not sufficient to claim the model can generalise to any unseen data. It is used to ensure that the model does not ‘over-fit’ to the training data. This is then repeated until a point defined by the user, usually either when a certain number of epochs, or iterations, have been performed or when the loss has gone below a pre-determined threshold.
Some NNs make use of an encoder and a decoder within the code. The encoder block takes the input data and abstracts it into a latent vector, similar to the first two layers in Figure 1, whereas the decoder takes the latent vector and aims to reconstruct the input as the output. In this paper, there is an encoder before each model and a decoder after it. They are both defined using MLP and are trained along with the model.
@@
The dataset being used to compare the benchmarks of windAI_bench [1] to is AirfRANS, published by Extrality [2], due to its pre-existing benchmark of fluid flow over an aerofoil. It includes 1,000 aerofoils that have been benchmarked using four architectures: a basic MLP, PointNet [20], GraphSAGE [21], and GUNet [22], which is a GNN-based variant of the UNet architecture [23]. As mentioned in the introduction, further insight is given into the two datasets in section 3.1.
@@ 3 Methodology
The code for this paper was done using PyTorch and PyTorch Geometric [24] and will be made publicly available after acceptance.
@@ 3.1 Data
WindAI_bench includes two different models, transition and turbulent, both run by a 2D Hamiltonian Strand Navier-Stokes solver (HAM2D). Specifically being used is the airfoil_2k subsection of the dataset, which contains 1830 unique aerofoil shapes at 25 angles of attack, ranging from $-4^{\circ}$ to $20^{\circ}$, for 3 separate Reynolds numbers. This gives a total of $45,750$ simulations to train with, making the dataset comparable to that used by [25]. In order to give an insight into the amount of data required for reasonable results to be achieved by the models, only a limited number of aerofoils (across every angle of attack (AoA)) was utilised within this paper. Four sets of simulations was performed with 5, 20, 55 and 150 aerofoils for each of the models, the number of foils for each set has been chosen in line with an exponential scale, such that: $5\approx e^{1.5}$, $20\approx e^{3}$, $55\approx e^{4}$ and $150\approx e^{5}$. Whilst a full breakdown of the data can be found on windAI_bench’s GitHub [1], for the purposes of this paper, the data used was the $Re=3e^{6}$ data generated by the turbulence model. This includes, for each AoA, the $C_{L}$ and $C_{D}$ values, the ‘landmarks’ array which is a coordinate file for each point along the aerofoil, and the physical parameters at each node: namely the $x$ and $y$ positional co-ordinates, the density of the fluid, the $x$ and $y$ components of the momentum, the total energy, and the vorticity.
@@
The dataset was also normalised, making it non-dimensional as this reduces any potential inaccuracies caused by performing mathematical operations on largely differing values. The normalised values are found as in Eq. 2.
@@
The errors on the nodes farther away from the aerofoil were typically very low when compared to those close to the aerofoil, causing the overall error of the model to be artificially low, creating a large error at the points in which this paper is interested. Therefore, when loading the data, all points outside a radius of $0.7m$, centred at the midpoint of the chord length ($c=1m$), were omitted, noticeably increasing the error of the system, but ensuring better results at the points of interest. A graphical interpretation of this can be seen in Figure 2.
@@ 3.2 Models
The loss criterion used for the models during training is the Mean Squared Error (MSE) as outlined in Eq. 3.
@@ 3.3 Post Processing
To properly assess the effectiveness of the trained models, a validation check is performed using 60 unseen aerofoils to assess the ability of the models to generalise to relevant unseen data. The metrics of interest for this analysis are the root mean squared error (RMSE) of both the prediction and the $C_{L}$ calculated from the predicted flow. RMSE is more favourable than MSE as it converts the results back to the original scale and units of the target value, making it an easier metric to interpret. The formulation for the RMSE can be seen in Eq. 4.
@@
The panel method is a method that allows for the approximation of forces acting upon an object in a flow by segmenting that object, in this case an aerofoil, into discrete surface elements, panels, and prescribing a flow element to it [29]. Using the volumetric vorticity ($\omega$) predicted at each node on the aerofoil, the total circulation ($\Gamma$) for the aerofoil can be calculated by integrating $\omega$ over the total area defined by the aerofoil as seen in Eq. 5.
@@
This allows for the calculation of the Lift ($L$) and coefficient of lift $C_{L}$ acting on the aerofoil to be calculated as seen in Eqs. 6 & 7 [29]:
@@ 4 Results
As mentioned earlier in Section 3.2, MSE was used to calculate the loss during the training and testing cycle, as is common practice, after which the final losses were converted into RMSE and presented in the tables below, in order to provide a more insightful view into accuracy. The graphs below are laid out as follows: for each total number of aerofoils, the first table contains the RMSE during testing at points on the aerofoil surface (Surface) and across all other points that are not touching the aerofoil (Fluid). Then, the last three columns of the first table contain the results from the validation as laid out in Section 3.3, with the RMSE at the surface and in the fluid (the same points highlighted in the test RMSE) as well as the RMSE of the prediction of $C_{L}$ averaged over all 60 aerofoils (1,500 simulations). The second and third sections for each table contain a breakdown of the RMSE during testing of each of the variables being predicted, namely density ($\rho$), x-momentum ($\sigma_{u}$), y-momentum ($\sigma_{v}$), energy ($e$), and vorticity ($\omega$).
@@ 5 Aerofoils
Model Test RMSE Validation RMSE
@@
MLP 0.847 0.388 1.254 0.467 1.162
PointNet 1.302 0.484 1.643 0.603 0.953
GraphSAGE 0.750 0.390 2.299 2.920 1.219
GUNet 0.702 0.361 2.207 0.913 1.250
Validation RMSE Per Variable at Surface
@@
MLP 0.686 0.541 0.426 0.771 2.514
PointNet 0.644 1.117 0.906 0.748 3.233
GraphSAGE 1.399 0.742 0.753 1.618 4.554
GUNet 1.019 0.950 0.664 1.183 4.536
Validation RMSE Per Variable in Fluid
@@
MLP 0.432 0.489 0.375 0.431 0.581
PointNet 0.418 0.564 0.497 0.434 0.943
GraphSAGE 1.571 1.030 0.738 1.867 5.921
GUNet 0.627 0.859 0.654 0.653 1.478
@@ 20 Aerofoils
Model Test RMSE Validation RMSE
@@
MLP 0.422 0.234 0.640 0.263 0.933
PointNet 0.491 0.255 0.727 0.315 0.811
GraphSAGE 0.348 0.201 2.095 0.843 0.982
GUNet 0.422 0.247 1.917 0.949 1.290
Validation RMSE Per Variable at Surface
@@
MLP 0.432 0.150 0.097 0.477 1.266
PointNet 0.442 0.403 0.271 0.485 1.405
GraphSAGE 1.324 0.785 0.430 1.488 4.145
GUNet 1.247 0.532 0.369 1.405 3.799
Validation RMSE Per Variable in Fluid
@@
MLP 0.272 0.285 0.243 0.272 0.241
PointNet 0.284 0.335 0.286 0.288 0.373
GraphSAGE 0.745 0.675 0.703 0.746 1.222
GUNet 0.673 0.621 0.683 0.687 1.652
@@ 55 Aerofoils
Model Test RMSE Validation RMSE
@@
MLP 0.341 0.140 0.456 0.181 0.894
PointNet 0.351 0.148 0.431 0.178 0.817
GraphSAGE 0.248 0.109 1.934 0.778 2.045
GUNet 0.279 0.130 1.921 0.648 1.733
Validation RMSE Per Variable at Surface
@@
MLP 0.314 0.108 0.078 0.348 0.897
PointNet 0.261 0.139 0.109 0.288 0.863
GraphSAGE 1.266 0.522 0.326 1.413 3.836
GUNet 1.148 0.495 0.240 1.287 3.894
Validation RMSE Per Variable in Fluid
@@
MLP 0.190 0.198 0.159 0.190 0.165
PointNet 0.163 0.212 0.157 0.167 0.184
GraphSAGE 0.734 0.694 0.692 0.707 1.013
GUNet 0.553 0.571 0.556 0.547 0.927
@@
Model Test RMSE Validation RMSE
@@
MLP 0.204 0.024 0.355 0.134 0.890
PointNet 0.309 0.033 0.400 0.147 0.933
GraphSAGE 0.154 0.019 1.833 0.743 0.941
GUNet 0.215 0.026 2.251 0.510 1.264
Validation RMSE Per Variable at Surface
@@
MLP 0.235 0.084 0.055 0.259 0.705
PointNet 0.266 0.109 0.076 0.296 0.790
GraphSAGE 1.263 0.489 0.269 1.406 3.594
GUNet 0.741 0.636 1.222 0.920 4.694
Validation RMSE Per Variable in Fluid
@@
MLP 0.141 0.152 0.109 0.141 0.126
PointNet 0.153 0.160 0.124 0.152 0.143
GraphSAGE 0.749 0.650 0.623 0.745 0.914
GUNet 0.347 0.336 0.404 0.349 0.885
@@
Figure 4: Bar Chart showing the RMSE between the calculated and real $C_{L}$ for all 60 validation aerofoils

### v2

@@ Abstract
This paper investigates the capability of Neural Networks (NNs) as alternatives to the traditional methods to analyse the performance of aerofoils used in the wind and tidal energy industry. The current methods used to assess the characteristic lift and drag coefficients include Computational Fluid Dynamics (CFD), thin aerofoil and panel methods, all face trade-offs between computational speed and the accuracy of the results and as such NNs have been investigated as an alternative with the aim that it would perform both quickly and accurately. As such, this paper provides a benchmark for the windAI_bench dataset published by the National Renewable Energy Laboratory (NREL) in the USA. In order to validate the methodology of the benchmarking, the AirfRANS dataset benchmark is used as both a starting point and a point of comparison.
This study evaluates four neural networks (MLP, PointNet, GraphSAGE, GUNet) trained on a range of aerofoils at 25 angles of attack (4 to 20) to predict fluid flow and calculate lift coefficients ($C_{L}$) via the panel method. GraphSAGE and GUNet performed well during the training phase, but underperformed during testing. Accordingly, this paper has identified PointNet and MLP as the two strongest models tested, however whilst the results from MLP are more commonly correct for predicting the behaviour of the fluid, the results from PointNet provide the more accurate results for calculating $C_{L}$.
@@ 1 Introduction
In order to validate the methodology applied to benchmarking windAI_bench, the second dataset being investigated is the AirfRANS dataset published by Bonnet et al. [9]. AirfRANS contains 1,000 aerofoils and is benchmarked for use in Machine Learning (ML) by using four separate architectures. Importantly, whilst it includes training and validation loops, AirfRANS does not incorporate testing against unseen geometries or foils, meaning that the ability of any of the benchmark models to generalise to completely unseen data has not been verified. With this in mind, whilst AirfRANS will be of great use for a comparison of methodologies used for benchmarking this type of data, it can not be used for comparison of results due to the fundamental difference between the two. To keep the results in line with those for similar datasets, the models used will be the same as those in Bonnet et al. [9].
One point worth noting is that whilst these datasets are designed for wind turbines, they are also applicable to the design and evaluation of tidal turbines. To account for scalability issues, Reynolds number calculations would thus need to consider flow characteristics and operational velocities at which tidal turbines usually operate.
As mentioned, the computational requirement of generating CFD data means that the data is regularly unattainable in large enough quantities for data-driven NNs to be viable. This paper aims to investigate the amount of data required to attain reasonable results. Therefore, the performance of each model has been investigated over 5, 20, 55 and 150 aerofoils. These numbers were chosen to be roughly equal to an exponential scale as $5\approx e^{1.5}$, $20\approx e^{3}$, $55\approx e^{4}$ and $150\approx e^{5}$. It is important to validate the results from this against a sample set of aerofoils not used in the training of the model, i.e., a test dataset, to ensure that the model is not over-fitting to the data presented.
@@ 2 Related Literature
Artificial Neural Networks (ANNs), or simply NNs, are universal function approximators [10] used within ML for the prediction of specified functions. NNs are called so because they are based on the design of the biological neural networks found within a brain, with neurons and synapses represented by the neurons that make up each layer of the NN and the connections that connect them, a diagram of which can be seen in Fig. 1. The network is divided up into three categories of layers (input, hidden and output) each of a user specified width, i.e. how many neurons in each layer. Each connection contains a weight ($W$) which can be within the range $0\leq W\leq 1$. These are multiplied by the output from the neuron it comes from ($x$) before being summed up with the remaining connections along with a bias factor ($b$), as seen in (1).
@@
These weights and neuron values are used to affect the final output of the NN and this output is then compared to a ‘ground truth’ value, i.e. the ‘real’ value usually taken from the data. Then, after finding the gradient of the ‘loss’, or difference between these two values, the weights and biases of the network are updated using an optimisation algorithm such as ADAM [11]. The data loaded into the model is split into training and validation data: the former is used to optimise the weights and biases of the model: whilst the latter is used to check the ability of the model to generalise to unseen data. It is worth noting that incorporating a validation split of data is not sufficient to claim the model can generalise to any unseen data. It is used to ensure that the model does not ‘over-fit’ to the training data. This is then repeated until a point defined by the user, usually either when a certain number of epochs, or iterations, have been performed or when the loss has gone below a pre-determined threshold.
Some NNs make use of an encoder and a decoder within the code. The encoder block takes the input data and abstracts it into a latent vector, similar to the first two layers in Fig. 1, whereas the decoder takes the latent vector and aims to reconstruct the input as the output. In this paper, there is an encoder before each model and a decoder after it. They are both defined using MLP and are trained along with the model.
@@
The dataset being used to compare the benchmarks of windAI_bench [8] to is AirfRANS, published by Bonnet et al. [9], due to its pre-existing benchmark of fluid flow over an aerofoil. It includes 1,000 aerofoils that have been benchmarked using four architectures: a basic MLP, PointNet [20], GraphSAGE [21], and GUNet [22], which is a GNN-based variant of the UNet architecture [23]. As mentioned in the introduction, further insight is given into the two datasets in section 3.1.
@@ 3 Methodology
The code for this paper was done using PyTorch and PyTorch Geometric [24] and is available on GitHub: https://github.com/OllieS-PhD/Benchmark_Aerofoils
@@ 3.1 Data
WindAI_bench includes two different models, transition and turbulent, both run by a 2D Hamiltonian Strand Navier-Stokes solver (HAM2D). Specifically being used is the airfoil_2k subsection of the dataset, which contains 1830 unique aerofoil shapes at 25 angles of attack, ranging from $-4^{\circ}$ to $20^{\circ}$, for 3 separate Reynolds numbers. This gives a total of $45,750$ simulations to train with, making the dataset comparable to that used by [25]. In order to give an insight into the amount of data required for reasonable results to be achieved by the models, only a limited number of aerofoils (across every angle of attack (AoA)) were utilised within this paper. Four sets of simulations was performed with 5, 20, 55 and 150 aerofoils for each of the models, the number of foils for each set has been chosen in line with an exponential scale, such that: $5\approx e^{1.5}$, $20\approx e^{3}$, $55\approx e^{4}$ and $150\approx e^{5}$. Whilst a full breakdown of the data can be found on windAI_bench’s GitHub [8], for the purposes of this paper, the data used was the $Re=3e^{6}$ data generated by the turbulence model. This includes, for each AoA, the $C_{L}$ and $C_{D}$ values, the ‘landmarks’ array which is a coordinate file for each point along the aerofoil, and the physical parameters at each node: namely the $x$ and $y$ positional co-ordinates, the density of the fluid, the $x$ and $y$ components of the momentum, the total energy, and the vorticity.
@@
The dataset was also normalised, making it non-dimensional as this reduces any potential inaccuracies caused by performing mathematical operations on largely differing values. The normalised values are found as in (2).
@@
The errors on the nodes farther away from the aerofoil were typically very low when compared to those close to the aerofoil, causing the overall error of the model to be artificially low, creating a large error at the points in which this paper is interested. Therefore, when loading the data, all points outside a radius of $0.7m$, centred at the midpoint of the chord length ($c=1m$), were omitted, noticeably increasing the error of the system, but ensuring better results at the points of interest. A graphical interpretation of this can be seen in Fig. 2.
@@ 3.2 Models
The loss criterion used for the models during training is the Mean Squared Error (MSE) as outlined in (3).
@@ 3.3 Post Processing
To properly assess the effectiveness of the trained models, testing is performed using 60 unseen aerofoils to assess the ability of the models to generalise to relevant unseen data. The metrics of interest for this analysis are the root mean squared error (RMSE) of both the prediction and the $C_{L}$ calculated from the predicted flow. RMSE is more favourable than MSE as it converts the results back to the original scale and units of the target value, making it an easier metric to interpret. The formulation for the RMSE can be seen in (4).
@@
The panel method is a method that allows for the approximation of forces acting upon an object in a flow by segmenting that object, in this case an aerofoil, into discrete surface elements, panels, and prescribing a flow element to it [29]. Using the volumetric vorticity ($\omega$) predicted at each node on the aerofoil, the total circulation ($\Gamma$) for the aerofoil can be calculated by integrating $\omega$ over the total area defined by the aerofoil as seen in (5).
@@
This allows for the calculation of the Lift ($L$) and coefficient of lift $C_{L}$ acting on the aerofoil to be calculated as seen in 6 & 7 [29]:
@@ 4 Results
As mentioned earlier in Section 3.2, MSE was used to calculate the loss during the training cycle, as is common practice, after which the final losses were converted into RMSE and presented in the tables below, in order to provide a more insightful view into accuracy. The tables below are laid out as follows: for each total number of aerofoils, the first table contains the RMSE during training at points on the aerofoil surface (Surface) and across all other points that are not touching the aerofoil (Fluid). Then, the last three columns of the first table contain the results from testing as laid out in Section 3.3, with the RMSE at the surface and in the fluid (the same points highlighted in the test RMSE) as well as the RMSE of the prediction of $C_{L}$ averaged over all 60 aerofoils (1,500 simulations). The second and third sections for each table contain a breakdown of the RMSE during testing of each of the variables being predicted, namely density ($\rho$), x-momentum ($\sigma_{u}$), y-momentum ($\sigma_{v}$), energy ($e$), and vorticity ($\omega$).
@@ 5 Aerofoils
Model Train RMSE Test RMSE
@@
MLP 1.135 0.470 1.630 0.510 0.873
PointNet 1.298 0.448 1.703 0.483 1.019
GraphSAGE 0.718 0.342 2.157 1.109 1.020
GUNet 0.707 0.337 2.340 0.812 1.568
Test RMSE Per Variable at Surface
@@
MLP 0.716 0.785 0.816 0.812 3.292
PointNet 0.690 0.914 0.816 0.784 3.452
GraphSAGE 1.210 0.802 0.868 1.406 4.294
GUNet 1.313 1.221 0.641 1.522 4.629
Test RMSE Per Variable in Fluid
@@
MLP 0.460 0.483 0.418 0.463 0.684
PointNet 0.429 0.473 0.381 0.429 0.657
GraphSAGE 0.793 0.963 0.657 0.845 1.857
GUNet 0.745 0.859 0.679 0.742 0.996
@@ 20 Aerofoils
Model Train RMSE Test RMSE
@@
MLP 0.587 0.297 0.788 0.311 0.997
PointNet 0.747 0.396 0.881 0.358 1.015
GraphSAGE 0.563 0.288 1.820 1.024 1.394
GUNet 0.578 0.306 1.808 0.653 1.335
Test RMSE Per Variable at Surface
@@
MLP 0.599 0.222 0.124 0.658 1.498
PointNet 0.602 0.470 0.249 0.664 1.673
GraphSAGE 1.262 0.599 0.472 1.404 3.523
GUNet 1.097 0.625 0.454 1.249 3.604
Test RMSE Per Variable in Fluid
@@
MLP 0.359 0.317 0.243 0.352 0.267
PointNet 0.354 0.365 0.271 0.353 0.429
GraphSAGE 0.750 0.741 0.720 0.769 1.737
GUNet 0.633 0.638 0.515 0.628 0.817
@@ 55 Aerofoils
Model Train RMSE Test RMSE
@@
MLP 0.560 0.217 0.729 0.261 0.795
PointNet 0.638 0.243 0.764 0.268 0.748
GraphSAGE 0.472 0.191 1.543 0.760 1.299
GUNet 0.617 0.228 1.830 0.615 1.493
Test RMSE Per Variable at Surface
@@
MLP 0.551 0.115 0.072 0.610 1.402
PointNet 0.544 0.157 0.111 0.602 1.492
GraphSAGE 1.011 0.397 0.602 1.129 3.015
GUNet 1.161 0.380 0.328 1.311 3.663
Test RMSE Per Variable in Fluid
@@
MLP 0.310 0.247 0.198 0.304 0.228
PointNet 0.308 0.265 0.209 0.301 0.243
GraphSAGE 0.617 0.712 0.736 0.605 1.046
GUNet 0.624 0.530 0.519 0.606 0.763
@@
Model Train RMSE Test RMSE
@@
MLP 0.428 0.146 0.335 0.138 0.861
PointNet 0.586 0.186 0.458 0.171 0.954
GraphSAGE 0.472 0.153 1.946 0.913 1.168
GUNet 0.536 0.173 2.384 0.600 1.358
Test RMSE Per Variable at Surface
@@
MLP 0.212 0.107 0.061 0.233 0.668
PointNet 0.284 0.158 0.102 0.315 0.913
GraphSAGE 1.318 0.292 0.313 1.467 3.856
GUNet 1.006 0.377 0.760 1.222 5.019
Test RMSE Per Variable in Fluid
@@
MLP 0.136 0.159 0.121 0.137 0.131
PointNet 0.172 0.183 0.144 0.172 0.181
GraphSAGE 0.733 0.594 0.600 0.725 1.547
GUNet 0.444 0.436 0.401 0.447 1.025
@@
Figure 4: Bar Chart showing the RMSE between the calculated and real $C_{L}$, averaged over every angle of attack for all 60 validation foils